# Compilation for Dynamically Field-Programmable Qubit Arrays with Efficient and Provably Near-Optimal Scheduling

### v1

@@
Abstract
Dynamically field-programmable qubit arrays based on neutral atoms have high fidelity and highly parallel gates for quantum computing. However, it is challenging for compilers to fully leverage the novel flexibility offered by such hardware while respecting its various constraints. In this study, we break down the compilation for this architecture into three tasks: scheduling, placement, and routing. We formulate these three problems and present efficient solutions to them. Notably, our scheduling based on graph edge coloring is provably near-optimal in terms of two-qubit gate stage count (at most one more than the optimum), the fidelity bottleneck of this platform. As a result, our compiler, Enola, produces higher fidelity results compared to existing works, e.g., 3.7X stage reduction and 5.9X fidelity improvement on the benchmark set used by OLSQ-DPQA, the current state of the art. Additionally, Enola is highly scalable, e.g., within 30 minutes, it can compile circuits with 10,000 qubits, a scale sufficient for the current era of quantum computing. Enola is open source at https://github.com/UCLA-VAST/Enola.
I Introduction
In recent years, quantum computing based on neutral atoms has advanced quickly in scale, quality, and adoption. The largest experiment so far captured 6,100 qubits [1], greater than other quantum computing platforms. One-qubit gates with 99.97% fidelity and two-qubit gates with 99.5% fidelity have been demonstrated [2, 3] to be competitive among the platforms. As a result, in addition to intensifying academic efforts, multiple startup companies [4, 5, 6, 7, 8] have been established to pursue this route of quantum computing.
A particular advantage of neutral atoms is the ability to move the qubits. Via these movements, the coupling among qubits can be field-programmable dynamically in different stages of the quantum circuit execution. This allows for a lot more flexibility to apply two-qubit entangling gates that are essential to quantum computing. Thus, researchers were able to run some of the most advanced quantum circuits requiring non-local connectivity on the dynamically field-programmable qubit arrays (DPQA) architecture [2, 9].
@@ I Introduction
Figure 1: Dynamically field-programmable qubit arrays (DPQA). a) Qubits (blue dots) can transfer between SLM traps (circles) and AOD traps (intersections of red lines). AOD rows and columns can move while preserving their relative order. b) A global Rydberg laser excites all qubits. A two-qubit gate is applied if two qubits are within the Rydberg range.
Figure 2: OLSQ-DPQA and Enola error breakdown. The benchmarks are 3-regular MaxCut QAOA circuits used in Ref. [10]. For the 90-qubit circuits, Elona reduces two-qubit gate stages by 3.7X and improves the overall fidelity by 5.9X.
@@
The compilation for DPQA involves a few tasks: scheduling which assigns two-qubit gates to Rydberg stages, placement which maps the qubits to sites initially, and routing which transfers and moves qubits between stages. The qubit mapping [11, 12, 13, 14, 15, 16, 17, 18] or quantum layout synthesis (QLS) problem [19, 20, 21, 22, 23, 24] ties these tasks together. OLSQ-DPQA [25, 10] is the first work to investigate QLS for DPQA with an effort to find the optimal QLS solution. Subsequent works Q-Pilot [26] and Atomique [27] aim to develop scalable heuristic solutions to support solving large-scale QLS problems for DPQA. However, we find opportunities for improvement both in quality and scalability. Specifically, the formulation of OLSQ-DPQA hinders the exploration of multiple rearrangement steps between Rydberg stages, resulting in more stages than the theoretical lower bound. Through a fidelity estimation detailed later, we find the two-qubit gates (including the side effect errors) to be the dominating error source in their results, as seen in Fig. 2a. In terms of scalability, OLSQ-DPQA can only handle circuits with up to 90 qubits in a day because it depends on solving satisfiability modulo theories (SMT) problems, which are NP-complete.
In this work, we formulate the three tasks in QLS individually and present efficient solutions for them. As a result, our compiler Enola (efficient and near-optimal layout synthesizer for atom arrays) has both higher quality and scalability than OLSQ-DPQA. The quality improvements of Enola are mainly due to the reduction of Rydberg stages, as indicated by the suppression of the two-qubit gate portion in Fig. 2b compared to Fig. 2a. Specifically, we can model two-qubit gates as edges in a graph so assigning gates to stages becomes coloring the edges in the graph. Suppose the optimal number of Rydberg stages is $S_{\text{opt}}$. Leveraging an efficient and provably near-optimal edge-coloring algorithm [28], Enola manages to schedule the gates to $S_{\text{opt}}$ or $S_{\text{opt}}+1$ stages. The placement problem is solved by simulated annealing to reduce the qubit traveling distance, and the routing solution is generated by solving independent sets to avoid AOD order violations. For the 90-qubit QAOA 3-regular MaxCut benchmarks [10], Enola produces 3.7X fewer stages and improves the overall fidelity by 5.9X. Furthermore, Enola can handle much larger circuits because it consists of scalable algorithms. We demonstrate compiling circuits with up to 10,000 qubits in 30 minutes, compared to OLSQ-DPQA’s 90 qubits in a day.
@@
The paper is organized as follows. Sec. II motivates the paper by analyzing the OLSQ-DPQA results with a detailed fidelity model. The following three sections (III, IV, and V) provide our solutions of scheduling, placement, and routing. Next, Sec. VI presents the evaluations. Then, Sec. VII introduces related works. Finally, we conclude the paper and suggest future directions in Sec. VIII. Additionally, we attach an appendix on the NP-hardness of optimal routing.
II Motivation: Fidelity Analysis
In OLSQ-DPQA, the SMT model only gives a high-level description: the interaction site where each qubit is at for each Rydberg stage. Given this SMT solution, a software module called CodeGen derives the other details like when and where to turn on or off the AOD. These details are enclosed in the generated DPQA code file consisting of a small set of DPQA instructions: initialization, Rydberg laser exposure, turning on/off the AOD, and AOD movements. Ref. [10] mainly reports the number of Rydberg stages, but with the CodeGen infrastructure, we can examine their compiled results with a more careful fidelity model.
@@ II Motivation: Fidelity Analysis
$\begin{split}f\ =\ (f_{1})^{g_{1}}\ &\cdot\ \overbrace{(f_{2})^{g_{2}}\ \cdot\ (f_{\text{exc}})^{|Q|S-2g_{2}}}^{\text{two-qubit gate}}\ \cdot\ \overbrace{(f_{\text{trans}})^{N_{\text{trans}}}}^{\text{atom transfer}}\\ &\quad\quad\quad\cdot\ \underbrace{\prod_{q\in Q}\ (1-T_{q}/T_{2})}_{\text{decoherence}}\ ,\end{split}$ (1)
@@
In Ref. [10], OLSQ-DPQA compiles a set of QAOA circuits designed for the MaxCut problem on 3-regular graphs [30] with the number of qubits ranging from 30 to 90. The specific circuit is the problem unitary in QAOA, $U_{C}$, consisting of 3 commutable two-qubit gates on each qubit. We evaluate the compiled results with our fidelity model and present the breakdown in Fig. 2a. Note that, to draw the figure, we take the logarithm of the fidelity terms so that they are additive. At 90 qubits, the two-qubit gate fidelity term is 0.0414, the atom transfer term is 0.592, and the decoherence term is 0.223. Thus, the dominating error source are the two-qubit gates, confirming what the authors claim in Ref. [10]. However, there is a gap between the number of stages achieved by OLSQ-DPQA, on average 14.6 for 90 qubits, and the theoretical lower bound, 3, because each qubit is only involved in 3 two-qubit gates. With our compiler, Enola, only 4 stages are produced, pushing the two-qubit fidelity term to 0.406. This effect is evident in the great decrease of the two-qubit gate portion in Fig. 2b compared to Fig. 2a.
III Scheduling: Edge Coloring
Figure 3: Scheduling in Enola. a) Scheduling a commutation group of two-qubit gates with edge coloring. b) Generic circuits can be considered as dependency subcircuits and commutation groups. Gates in the former is scheduled ASAP.
The two-qubit gates available on DPQA are controlled rotation in the Z basis, which are known to commute [31]. This means a set of these two-qubit gates can be executed in any order. We can solve the scheduling of a commutation group, i.e., a set of commutable two-qubit gates, with graph theoretic tools.
For a graph $G=(V,E)$, an edge coloring is a function $\phi:E\to\mathbb{Z}$ that evaluates different values for two different edges incident on a vertex, i.e., for $e,e^{\prime}\in E$ and $e\neq e^{\prime}$, if $e\cap e^{\prime}\neq\emptyset$, then $\phi(e)\neq\phi(e^{\prime})$. Misra&Gries [28] provide an algorithm with runtime $O(|V|\cdot|E|)$ that gives an edge coloring $\Phi:E\to\{0,1,...,\Delta(G)\}$ where $\Delta(G)$ is the maximum vertex degree of $G$. Thus, $\Phi$ colors the edges with at most $\Delta(G)+1$ colors. The minimum number of colors to color the edges is called the chromatic index of the graph, $\chi^{\prime}(G)$. Since $\chi^{\prime}(G)\geq\Delta(G)$, $\Phi$ colors the edges with at most $\chi^{\prime}(G)+1$ colors. Our compiler leverages these results to schedule commutable two-qubit gates. Its advantage is due to the following theorem.
@@ III Scheduling: Edge Coloring
Proof. A commutation group of two-qubit gates can be represented by a qubit interaction graph $G=(V,E)$ where the vertices are qubits and the edges are the two-qubit gates (Fig. 3a). The schedule is a function $\psi:E\to\mathbb{N}$ such that a qubit can only be involved in one gate at a Rydberg stage, i.e., for $e,e^{\prime}\in E$ and $e\neq e^{\prime}$, if $\psi(e)=\psi(e^{\prime})$, then $e\cap e^{\prime}=\emptyset$. This is contrapositive to the definition of an edge coloring, so $\psi$ is an edge coloring. Thus, the optimal number of Rydberg stages is $S_{\text{opt}}=\chi^{\prime}(G)$, which means the function $\Phi$ derived by the Misra&Gries algorithm maps the two-qubit gates to at most $S_{\text{opt}}+1$ Rydberg stages. Since $|E|$ is $O(n^{2})$ where $n$ is the number of qubits, and the Misra&Gries algorithm is $O(|V|\cdot|E|)$, the time complexity of our scheduling is $O(n^{3})$. ∎
@@
A more generic quantum circuit is specified by a sequence of gates. If two gates act on the same qubit, their relative order dictates a dependency. In Fig. 3b, we exhibit an example of of how one derives the dependency DAG for the two-qubit gates in a generic circuit. In this case, the scheduling problem is straightforward: the optimal number of stages is the critical path in the DAG and ASAP scheduling can achieve optimality. Although there is a way to augment the DAG to represent partially commutable circuits [32], supporting this in general requires mixing logic synthesis and layout synthesis. Therefore, we make an assumption similar to Ref. [33] that the whole quantum circuit is sliced into subcircuits that either respect all derived dependencies, as ‘dependency subcircuits’ shown in Fig. 3b, or are commutation groups. The scheduling for the slices can be performed simultaneously and the results can be stitched together afterwards. This sliced structure is prevalent in quantum computing. An example is the graph state preparation with various applications [34], which has a layer of Hadamard gates in the beginning and then commuting CZ gates. Another example is MaxCut QAOA that has alternating driver unitaries $U_{B}$ with dependency and problem unitaries $U_{C}$ that are commutation groups of ZZ gates.
IV Placement: Simulated Annealing
Figure 4: Placement in Enola. a) Trivial placement from left to right, from top to bottom. b) Placement with gate distance optimized by simulated annealing. c) Dynamic placement: after a Rydberg stage (red) is executed (left), run simulated annealing on moved qubits for a new placement (right).
@@ IV Placement: Simulated Annealing
In placement, we map qubits to interaction sites. The two-qubit gates at each Rydberg stage thus correspond to 2-pin nets between the sites. If the nets have a long wire-length, it takes more time to move the qubits, resulting in more decoherence, the second largest error source. As an example, qubits can be placed trivially from left to right and from top to bottom as in Fig. 4a. Then, the total distance of gates in the commutation group in Fig. 3a accumulates to $25.67\times 2.5r_{b}$. In comparison, an optimized placement displayed in Fig. 4b achieves a total wire-length of $19.48\times 2.5r_{b}$. To minimize the qubit travelling distances, our cost function is defined as
$\sum\nolimits_{g(q,q^{\prime})\in G}\ w_{g}\ \cdot\ \text{dist}(m(q),m(q^{\prime})),$ (2)
@@
where $w_{g}$ is the weight for gate $g$, $m$ is the placement function from qubits to interaction sites, and ‘dist’ is the Euclidean distance.
@@
We apply a simulated annealing algorithm, Fast-SA [35], to optimize this cost due to its effectiveness in solving discrete optimization problems. To enhance exploration efficiency, we confine qubits to a specific region, thus reducing the search space. Assuming the number of qubits to place is $n$, and the interaction sites have column indices $\{0,1,...,x_{\text{max}}\}$ and row indices $\{0,1,...,y_{\text{max}}\}$, we define the chip region for exploration as $x\in[0,\max(\lfloor\sqrt{n}\rfloor+4,x_{\text{max}})]$ and $y\in[0,\max(\lfloor\sqrt{n}\rfloor+4,y_{\text{max}})]$. In Fig. 4, $x_{\text{max}}=3$ and $y_{\text{max}}=2$. Fast-SA has a three-stage annealing schedule to facilitate state space exploration. At the first stage, the temperature is high. In other words, we have higher probability to accept an inferior solution. This stage mimics a random search to explore a large solution space. Then, the second stage performs the pseudo-greedy local search with low temperature. The last stage is the hill-climbing search stage where the temperature increases again to escape from local minima. The state in the annealing process is a placement which we initialize randomly. Then, state transitions can be made by either reassigning a qubit to an empty site or exchanging the locations of two qubits. The annealing process will terminate if the temperature is lower than a threshold or the number of iterations exceed a predefined limit, so the placement algorithm has a constant runtime.
@@
The configuration after the first Rydberg stage (red) is on the left of Fig. 4c. The arrows indicate AOD movements from 4b to this configuration. At this point, we can always reverse the movements to return to Fig. 4b, and then find out the movements for the next stage (black). In this case, the placement is static for all the Rydberg stages so we set all the gate weights to 1 in the cost function.
@@
However, one can also consider dynamically changing the placement for the next stage. On the right of Fig. 4c, we display a new placement where the gate between $q_{5}$ and $q_{7}$ is shorter compared to Fig. 4b. If the placement is dynamic, gates earlier in the schedule should contribute more to the cost function. Thus, we set $w_{g}=\max(0.1,1-0.1s_{g})$, where $s_{g}$ is the number of stages preceding the stage that the gate $g$ belongs to, e.g., the gates in stages 0 to 3 will have the weights of 1, 0.9, 0.8, and 0.7, respectively. During the simulated annealing for intermediate placement, only the set of qubits necessitating relocation to vacant sites can be moved, while the remaining qubits must stay where they are. In our example, the new placement is restricted, from qubits $q_{2}$, $q_{5}$, $q_{7}$, and $q_{9}$ to the 6 empty sites. Placement of the other qubits are inherited from the previous placement. In our evaluations detailed later, dynamic placement slightly outperforms static placement. In a commutation group, there are at most $O(n)$ stages, so our placement runtime is $O(n)$, given that each placement spends a constant time.
V Routing: Independent Set

### v2

@@
Abstract.
Dynamically field-programmable qubit arrays based on neutral atoms feature high fidelity and highly parallel gates for quantum computing. However, it is challenging for compilers to fully leverage the novel flexibility offered by such hardware while respecting its various constraints. In this study, we break down the compilation for this architecture into three tasks: scheduling, placement, and routing. We formulate these three problems and present efficient solutions to them. Notably, our scheduling based on graph edge-coloring is provably near-optimal in terms of the number of two-qubit gate stages (at most one more than the optimum). As a result, our compiler, Enola, reduces this number of stages by 3.7x and improves the fidelity by 5.9x compared to OLSQ-DPQA, the current state of the art. Additionally, Enola is highly scalable, e.g., within 30 minutes, it can compile circuits with 10,000 qubits, a scale sufficient for the current era of quantum computing. Enola is open source at https://github.com/UCLA-VAST/Enola
journalyear: 2025copyright: rightsretainedconference: 30th Asia and South Pacific Design Automation Conference; January 20–23, 2025; Tokyo, Japanbooktitle: 30th Asia and South Pacific Design Automation Conference (ASPDAC ’25), January 20–23, 2025, Tokyo, Japandoi: 10.1145/3658617.3697778isbn: 979-8-4007-0635-6/25/01
1. Introduction
In recent years, quantum computing based on neutral atoms has advanced quickly in scale, quality, and adoption. Large experiments exceed 1,000 qubits (Pause et al., 2024), at the forefront of quantum computing. One-qubit gates with 99.97% fidelity and two-qubit gates with 99.5% fidelity have been demonstrated (Bluvstein et al., 2022; Evered et al., 2023) to be competitive among the platforms. As a result, in addition to intensifying academic efforts, multiple startup companies (QuEra, 2018; Infleqtion, 2007; Pasqal, 2019; planqc, 2022; Atom Computing, 2018) have been established to pursue this route of quantum computing.
A particular advantage of neutral atoms is the ability to move the qubits. Via these movements, the coupling among qubits is field-programmable dynamically in different stages of the quantum circuit execution. This allows for a lot more flexibility to apply two-qubit entangling gates that are essential to quantum computing. Thus, researchers were able to run some of the most advanced quantum circuits requiring non-local connectivity on the dynamically field-programmable qubit arrays (DPQA) architecture (Bluvstein et al., 2022, 2024).
@@ 1. Introduction
Figure 1. Dynamically field-programmable qubit arrays (DPQA). a) Qubits (blue dots) can transfer between SLM traps (circles) and AOD traps (intersections of red lines). AOD rows and columns can move while preserving their relative order. b) A global Rydberg laser excites all qubits. A two-qubit gate is applied if two qubits are within the Rydberg range.
Figure 2. Error breakdown of the OLSQ-DPQA and Enola results. The benchmarks are 3-regular MaxCut QAOA circuits used in Ref. (Tan et al., 2024a). For the 90-qubit circuits, Enola reduces two-qubit gate stages by 3.7x and improves the overall fidelity by 5.9x.
@@
Assuming the quantum circuit to execute has been decomposed into a proper gate set supported by DPQA, the compilation then involves a few tasks: scheduling gates to stages, placement of the qubits, and routing qubits between stages. For fixed architectures, these three tasks are often formulated together as the circuit placement (Maslov et al., 2008; Fan et al., 2022), qubit mapping (Li et al., 2019; Wille et al., 2019; Tan and Cong, 2021a; Huang et al., 2022; Li et al., 2023a; Park et al., 2022; Zhang et al., 2021; Molavi et al., 2022), or quantum layout synthesis (QLS) problem (Tan and Cong, 2021b, 2020; Lin et al., 2023; Wu et al., 2022; Shaik and van de Pol, 2023). OLSQ-DPQA (Tan et al., 2022, 2024a) is the first work to investigate QLS for DPQA with an effort to find the optimal QLS solutions. However, it can only handle circuits with up to 90 qubits in a day because it depends on solving satisfiability modulo theories (SMT) problems, which is NP-complete. The long compilation time hinders the adoption of this approach. Subsequent works Q-Pilot (Wang et al., 2024b) and Atomique (Wang et al., 2024a) aim to develop scalable heuristic solutions for better scalability. However, as indicated by the authors of these heuristic works themselves, the result quality is notably worse than OLSQ-DPQA.
In this work, we present a compiler Enola (efficient and near-optimal layout synthesizer for atom arrays) that is scalable and significantly improves result fidelity of previous works. In Enola, a scheduler first assigns quantum gates to stages. Then, a placer decides the location of qubits in each stage. Finally, a router derives the detail movement instructions between stages. On the 90-qubit 3-regular MaxCut QAOA (quantum approximate optimization algorithm) benchmarks (Tan et al., 2024a), Enola reduces the number of stage by 3.7x and improves the overall fidelity by 5.9x compared to OLSQ-DPQA. In terms of scalability, we demonstrate compiling circuits with up to 10,000 qubits in 30 minutes, which is sufficient for all DPQA hardware available now or in the near future.
The fidelity improvements of Enola are mainly due to the reduction of Rydberg stages, as indicated by the suppression of the two-qubit gate portion in Fig. 2b compared to Fig. 2a. Specifically, we can model two-qubit gates as edges in a graph so that assigning gates to stages becomes coloring the edges in the graph. Suppose the optimal number of Rydberg stages is $S_{\text{opt}}$. Leveraging the efficient and provably near-optimal Misra-Gries edge-coloring algorithm (Misra and Gries, 1992), Enola manages to schedule the gates to $S_{\text{opt}}$ or $S_{\text{opt}}+1$ stages. To our knowledge, we are the first to relate the DPQA scheduling problem to graph edge-coloring and leverage this near-optimal algorithm. In comparison, the formulation of OLSQ-DPQA hinders the exploration of multiple rearrangement steps between two stages, resulting in much more stages than the optimum.
In Enola, the placement problem is solved by simulated annealing to reduce the qubit traveling distance, and the routing solution is generated by independent sets to avoid AOD order violations while facilitating parallelism. To our knowledge, we are also the first to formulate the DPQA routing problem to solving independent sets.
@@
The paper is organized as follows. Sec. 2 motivates the paper by analyzing the OLSQ-DPQA results with a detailed fidelity model. The following three sections (3, 4, and 5) provide our solutions of scheduling, placement, and routing. Next, Sec. 6 presents the evaluations. Then, Sec. 7 introduces related works. Finally, in Sec. 8, we conclude the paper and suggest directions for future improvements that could address some limitations of this work.
2. Motivation: Fidelity Analysis
@@
(1) $f=(f_{1})^{g_{1}}\cdot\underbrace{(f_{2})^{g_{2}}\cdot(f_{\text{exc}})^{|Q|S-2g_{2}}}_{\text{two-qubit gate}}\cdot\underbrace{(f_{\text{trans}})^{N_{\text{trans}}}}_{\text{atom transfer}}\cdot\underbrace{\Pi_{q\in Q}\ (1-T_{q}/T_{2})}_{\text{decoherence}},$
@@
In Ref. (Tan et al., 2024a), OLSQ-DPQA compiles a set of QAOA circuits designed for the MaxCut problem on 3-regular graphs (Farhi et al., 2014) with the number of qubits ranging from 30 to 90. The specific circuit is the problem unitary in QAOA, $U_{C}$, consisting of 3 commutable two-qubit gates on each qubit. We evaluate their compiled results with our fidelity model and present the breakdown in Fig. 2a. Note that, to draw the figure, we take the logarithm of the fidelity terms so that they are additive. At 90 qubits, the two-qubit gate fidelity term is 0.0414, the atom transfer term is 0.592, and the decoherence term is 0.223. Thus, the dominating error source are the two-qubit gates, confirming claims Ref. (Tan et al., 2024a). However, there is a gap between their number of stages, on average 14.6 for 90 qubits, and the theoretical lower bound, 3, because each qubit is only involved in 3 two-qubit gates. Our compiler, Enola, produces only 4 stages, improving the two-qubit fidelity term to 0.406. This effect is evident in the great decrease of the two-qubit gate portion in Fig. 2b compared to Fig. 2a.
3. Scheduling: Edge Coloring
Figure 3. Scheduling in Enola. a) Scheduling a commutation group of two-qubit gates with edge coloring. b) Generic circuits can be divided to dependency subcircuits and commutation groups. Dependency subcircuits are scheduled ASAP.
Since the fidelity term of two-qubit gate is more critical than atom transfer and decoherence, and previous methods are far from reaching the optimal number of stages, our metric in scheduling is the number of stages. We first consider the scheduling of a special class of circuit, commutation groups, consisting of commutable two-qubit gates that can be executed in any order. A commutation group can be represented by a qubit interaction graph $G=(V,E)$ where the vertices are qubits and the edges are the two-qubit gates (Fig. 3a). The scheduling involves finding a function $\psi:E\to\mathbb{N}$ such that a qubit can only be involved in one gate at a Rydberg stage, i.e., for $e,e^{\prime}\in E$ and $e\neq e^{\prime}$, if $\psi(e)=\psi(e^{\prime})$, then $e\cap e^{\prime}=\emptyset$.
We now review some graph theory results. For a graph $G=(V,E)$, an edge coloring is a function $\phi:E\to\mathbb{Z}$ that evaluates different values for two different edges incident on a vertex, i.e., for $e,e^{\prime}\in E$ and $e\neq e^{\prime}$, if $e\cap e^{\prime}\neq\emptyset$, then $\phi(e)\neq\phi(e^{\prime})$. Misra-Gries (Misra and Gries, 1992) provide an algorithm with runtime $O(|V|\cdot|E|)$ that gives an edge coloring $\Phi:E\to\{0,1,...,\Delta(G)\}$ where $\Delta(G)$ is the maximum vertex degree of $G$. Thus, $\Phi$ colors the edges with at most $\Delta(G)+1$ colors. The minimum number of colors to color the edges is called the chromatic index of the graph, $\chi^{\prime}(G)$. Since $\chi^{\prime}(G)\geq\Delta(G)$, $\Phi$ colors the edges with at most $\chi^{\prime}(G)+1$ colors. Our compiler leverages these results to schedule commutable two-qubit gates. Its advantage is due to the following theorem.
@@ 3. Scheduling: Edge Coloring
Proof. Note that the above definition of a schedule function, $\psi$, is contrapositive to the definition of an edge coloring function, $\phi$. Thus, the scheduling function is just an edge coloring. Therefore, the optimal number of Rydberg stages $S_{\text{opt}}=\chi^{\prime}(G)$, which means the function $\Phi$ derived by the Misra-Gries algorithm maps the two-qubit gates to at most $S_{\text{opt}}+1$ Rydberg stages. Since $|E|$ is $O(n^{2})$ where $n$ is the number of qubits, and the Misra-Gries algorithm is $O(|V|\cdot|E|)$, the time complexity of our scheduling is $O(n^{3})$. ∎
@@
A more generic quantum circuit is specified by a sequence of gates. If two gates act on the same qubit, their relative order dictates a dependency. In Fig. 3b, we exhibit an example of how one derives the dependency DAG (directed acyclic graph) for the two-qubit gates in a generic circuit. In this case, the scheduling problem is straightforward: the optimal number of stages is the critical path in the DAG and ASAP (as soon as possible) scheduling can achieve optimality. Although there is a way to augment the DAG to represent partially commutable circuits (Iten et al., 2022), supporting this in general requires mixing logic synthesis and layout synthesis. Therefore, we make an assumption similar to Ref. (Shi et al., 2019) that the whole quantum circuit is sliced into subcircuits that either respect all derived dependencies, as ‘dependency subcircuits’ shown in Fig. 3b, or are commutation groups. The scheduling for the slices can be performed separately and the results can be stitched together afterwards. This sliced structure is prevalent in quantum computing. An example is the graph state preparation with various applications (Hein et al., 2006), which has a layer of Hadamard gates in the beginning and then commuting CZ gates. Another example is MaxCut QAOA that has alternating driver unitaries $U_{B}$ with dependency and problem unitaries $U_{C}$ that are commutation groups of ZZ gates.
In summary, for the dependency subcircuits, Enola uses ASAP scheduling which is optimal; for the commutation groups, Enola uses Misra-Gries algorithm, which is near-optimal. Thus, the scheduling in Enola is near-optimal in the number of stages.
4. Placement: Simulated Annealing
@@
In placement, we map qubits to interaction sites. The two-qubit gates at each Rydberg stage are like 2-pin nets in conventional circuit placement. If a net has a long wire-length, it takes more time to move the qubits, resulting in more decoherence, the second largest error source. Thus, we minimize the total wire-length in placement in order to reduce qubit movement time. Note that this is heuristic because some qubits movements can be simultaneous. As an example, qubits can be placed trivially from left to right and from top to bottom as in Fig. 4a. Then, the total distance of gates in the commutation group in Fig. 3a accumulates to $25.67\times 2.5r_{b}$. In comparison, an optimized placement displayed in Fig. 4b achieves a total wire-length of $19.48\times 2.5r_{b}$. To minimize the qubit traveling distances, our cost function is defined as
(2) $\sum\nolimits_{g(q,q^{\prime})\in G}\ w_{g}\cdot\text{dist}(\ m(q),m(q^{\prime})\ ),$
@@
where $w_{g}$ is the weight for gate $g$, $m$ is the circuit placement function from qubits to interaction sites, and ‘dist’ is the Euclidean distance. This is a common form of cost function in placement, and we apply a simulated annealing algorithm, Fast-SA (Chen and Chang, 2006), to optimize it.
Figure 4. Placement in Enola. a) Trivial placement from left to right, from top to bottom. b) Placement with gate distance optimized by simulated annealing. c) Dynamic placement: after a Rydberg stage (red) is executed (left), run simulated annealing on moved qubits for a new placement (right).
@@
To enhance exploration efficiency, we confine qubits to a specific region, thus reducing the search space. Assuming the number of qubits to place is $n$, and the interaction sites have column indices $\{0,1,...,x_{\text{max}}\}$ and row indices $\{0,1,...,y_{\text{max}}\}$, we define the chip region for exploration as $x\in[0,\max(\lfloor\sqrt{n}\rfloor+4,x_{\text{max}})]$ and $y\in[0,\max(\lfloor\sqrt{n}\rfloor+4,y_{\text{max}})]$. In Fig. 4, $x_{\text{max}}=3$ and $y_{\text{max}}=2$. Fast-SA has a three-stage annealing schedule to facilitate state space exploration. At the first stage, the temperature is high. In other words, we have higher probability to accept an inferior solution. This stage mimics a random search to explore a large solution space. Then, the second stage performs the pseudo-greedy local search with low temperature. The last stage is a hill-climbing search where the temperature increases again to escape from local minima. The state in the annealing process is a placement which we initialize randomly. Then, state transitions can be made by either reassigning a qubit to an empty site or exchanging the locations of two qubits. The annealing process will terminate if the temperature is lower than a threshold or the number of iterations exceed a predefined limit, so the placement algorithm has a constant runtime.
@@
The configuration after the first Rydberg stage (red) is on the left of Fig. 4c. The arrows indicate AOD movements from 4b to this configuration. At this point, we can always reverse the movements to return to Fig. 4b, and then find out the movements for the next stage (black). In this case, the placement is static for all the Rydberg stages, so we set all the gate weights to 1 in the cost function.
@@
However, one can also consider dynamically changing the placement for the next stage. On the right of Fig. 4c, we display a new placement where the gate between $q_{5}$ and $q_{7}$ is shorter compared to Fig. 4b. If the placement is dynamic, gates earlier in the schedule should contribute more to the cost function. Thus, we set $w_{g}=\max(0.1,1-0.1s_{g})$, where $s_{g}$ is the number of stages preceding the stage that the gate $g$ belongs to, e.g., the gates in stages 0 to 3 will have the weights of 1, 0.9, 0.8, and 0.7, respectively. During the simulated annealing for intermediate placement, we choose to move only the set of qubits necessitating relocation to vacant sites, while the remaining qubits stay where they are. In our example, the new placement is from qubits $q_{2}$, $q_{5}$, $q_{7}$, and $q_{9}$ to the 6 empty sites. The other qubits inherit the current placement. We observe that dynamic placement outperforms static placement, so Enola uses dynamic placement by default. In a commutation group, there are at most $O(n)$ stages, so our placement runtime is $O(n)$, given that each placement has a constant time limit.
5. Routing: Independent Set